A necklace purchased for Rs 1200 and sold at 35 percent loss. The second purchaser sells the same necklace for Rs 875 and pays 2 percent commission. find the profit or loss percent of the second purchaser.
Answer by akib_bd(2)

First seller sold it for Rs {1200-(1200x0.35)}= 780
The the second seller sold it for 875 and paid 2% commission, that means the money s/he get by selling the item is Rs {875-(875x.02)}=857.5.
Second seller bought it at Rs 780 and sold it at 857.5. So, s/he made at profit of Rs (857.5-780)=77.5 at a buying price of 780.
So s/he made a profit of (77.5/780)% =9.936%
:-)